1. (50pts) Figure out V and lo as the circuit shown below. Please show your works in details 2mA
Answer #1

ID Apply KVL for Loops Loop-192; super mesh Ik Ijtik (11-I3) + IK I2 +IKI2 +ık(Iz-1w) =0 Ist I 1-13 + 1 2 + I2 + I2 I4 =0 257from creuit Iz = 2 mA, I4 = 3 mA om from ea ll ① 2Ij+312 - Is-Iy = 0 21, +312 -2mA-3mA=0 211+3I2 = 5mA I solve call @ 85 I-I2